Write the expression sin 64 degrees in terms of cosine 

Grade:12

2 Answers

Karthik Vinod
7 years ago
sin (90-x)=cos (x) ( Trigonometry Indentity)
So, sin (90-26)= cos (26) = sin (64)
 
Cosine and Sine values appear to be related complementary, as in the above answer. If you’ve seen the pattern such as
cos (0) = sin (90)
cos (30) = sin (60)
cos (45) = sin(45)
cos(60) = sin (30)
cos (90) = sin (0),
The sum of the angles equal 90 degrees.
Karthik Vinod
7 years ago
I’m sorry I think I’ve misread your question. I suppose you’re asking sin 64 in terms of cos 64.
 
I think it will be
 
sin 64 = square root of (1 – cos2 64)
 
using sin2x + cos2 x =1 =